# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v1

@@ 5.1. Hypotheses
We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts scaffolding and affordances would provide participants a better experience. Our individual hypotheses were as follows:
@@
H4: Papers-to-Posts leads to lower task completion time than the baseline in generating a blog post that the writer would be comfortable posting publicly.
5.2. Participants
@@ 5.3. Study Conditions
The treatment condition involved interacting with Papers-to-Posts (Figure 2), which is described in Section 4. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections, so the resulting draft also had introduction, methods, results, and conclusion sections. However, the prompt did not contain section-specific instructions. The full prompt can be found in Appendix A.
@@
In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey. The paper was different for each condition. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The papers were parsed as described in Section 3.3. The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
5.4. Study Procedure
@@
The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline condition, in randomized order and following a counter-balanced design. The participants’ selected papers were assigned randomly to conditions. Each study session lasted up to about 75 minutes. To provide sufficient flexibility for recruited participants, they can select two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1). The condition order was counter-balanced within the group of participants with sessions 1 day apart and within the group of participants with sessions over 1 day apart.
@@
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Hangouts. To start, the session coordinator introduced participants to the task of writing a blog post with an AI tool. The coordinator then presented a tutorial on the tool. After, participants had 2 minutes to interact with the tool in the context of a sample paper that was not their own, and they could ask questions about the tool. Participants next opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area, which was computer science for all participants. We made this assumption because it was the most common audience for whom potential recruits wanted to write blog posts (52 of 60), according to the recruitment survey. The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any tools or information outside of the tool provided. We made an exception when participants wanted to view the PDF of their paper due to issues working with its parsing in the tool. This only occurred twice. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
@@
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, the coordinator reminded them to let them know if they reached the point of comfort with publication. If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added a rule so that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If and when the participant reached that point, or if an hour had passed, they were asked to submit and moved on to the follow-up survey regarding their experience with the tool. The survey consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals. If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions. These were not for analysis but only to help us get a quick sense of the results. However, they were not allowed to change their first-session responses. At the end of the session, the coordinator told participants to close out of the tool and, if it was their first session, not to use the tool before the next session. Each time the participant prompted the LLM, saved, or submitted, a log of the user action along with the full state of the blog post draft was saved for analysis.
5.5. Results
We evaluated our 4 hypotheses (Section 5.1) for statistical significance and found that blog post satisfaction is significantly different between the two conditions, while tool satisfaction, cognitive load, and difference in task completion time are not. We also investigated how well the design goals were fulfilled, and the extent of participants’ editing. We saw evidence that our tool addressed the design goals and increased participants’ satisfaction with control over the tool. In addition, we observed that participants did significantly more editing of the initial LLM draft using Papers-to-Posts (we note that this significance test was not part of our initial hypotheses and is therefore exploratory in nature).
5.5.1. Blog Post Satisfaction
We utilized the Likert responses (see Figure 5) as a measure of participants’ satisfaction with their final blog post under both conditions. Given the Likert nature of the data, we employed a non-parametric test. Furthermore, we used a sign test rather than a Wilcoxon signed rank test, which is normally used to analyze non-parametric within-subjects data for two conditions of one factor, due to the violation of that test’s assumption of a symmetric distribution of the paired differences about the median. We observed that participants were significantly more satisfied with the final blog post generated with Papers-to-Posts (M=6.00,Q1=5.00,Q3=7.00) as compared to the baseline tool (M=5.00,Q1=5.00,Q3=6.00) (Dependent-Samples Sign Test, S=11, p¡.05). The majority of participants (11) were more satisfied in the treatment condition, 7 were equally satisfied in both conditions, and only 2 were more satisfied in the baseline condition.
Inspecting participants’ Likert responses regarding aspects of their blog post satisfaction, we see that participants particularly demonstrated more satisfaction with the content covered in the Papers-to-Posts blog post than the baseline blog post (Figure 5b). The majority of participants (14) were more satisfied with the content under the treatment condition (equally: 2, less: 4). Overall, participants were more satisfied with the blog post structure using Papers-to-Posts than the baseline (Figure 5d). The plurality of participants (9) were more satisfied with the structure under the treatment (equally: 6, less: 5). Participants did not show increased satisfaction with the blog post’s tone under the treatment compared to the baseline (Figure 5c).
Looking at participants’ perceived task performance (Figure 6c), we observe that participants had higher perceived performance with Papers-to-Posts than with the baseline. The plurality of participants (9) thought they were more successful in writing a research-paper blog post with Papers-to-Posts compared to the baseline (equally: 7, less: 4). This further corroborates participants’ blog post satisfaction under the treatment condition.
5.5.2. Tool Satisfaction
For measuring tool satisfaction, we used the 5-question Likert scale on tool acceptance from Kocielnik et al. (Kocielnik et al., 2019) (Figure 5). The 5 responses were averaged to obtain one score. We used the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. While participants had a higher median satisfaction with Papers-to-Posts (M=5.9,Q1=5.2,Q3=6.5) than the baseline tool (M=5.6,Q1=4.8,Q3=6.0), this difference was not significant (Dependent-Samples Sign Test, S=13, p=n.s.).
In Figure 5e-h, we see that, under the treatment condition, participants were more likely to find the tool helpful, use it if it were available, recommend it, and perceive it as useful for productivity. For each of these questions, at least 4 more participants increased rather than decreased their Likert rating from the baseline to the treatment condition. However, for the question about the tool inducing annoyance, only 2 more participants indicated decreased rather than increased annoyance with Papers-to-Posts.
5.5.3. Cognitive Load
To measure cognitive load, we used the NASA TLX Index questionnaire (Hart and Staveland, 1988) (Figure 6a-d) and summed the responses for the mental demand, temporal demand, effort, and frustration questions (see 5.5.1 for discussion of the responses for performance). We utilized the non-parametric sign test to evaluate the data, as it was Likert data and violated the Wilcoxon signed rank test’s assumption of a symmetric distribution of the paired differences about the median. We saw no significant difference between the baseline and treatment conditions, although participants reported slightly lower cognitive load with Papers-to-Posts (M=13.0,Q1=11.5,Q3=16.3) than the baseline tool (M=14.5,Q1=12.0, Q3=18.3) (Dependent-Samples Sign Test, S=8, p=n.s.). Figure 6a-d presents little difference in mental demand, temporal demand, and frustration between the treatment and baseline conditions. However, with respect to effort, we note a divergence. The majority of participants (11) reported using less effort with Papers-to-Posts than with the baseline tool, whereas only 4 participants reported using more effort.
5.5.4. Task Completion Time
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test. Participants worked slightly faster with the baseline tool (M=44.3 minutes,SD=12.5) in comparison to Papers-to-Posts (M=46.0 minutes,SD=13.9), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
5.5.5. Control and Ownership
Participants indicated higher perceived control (treatment: M=5.00,Q1=4.50,Q3=6.00, baseline: M=5.00, Q1=3.00,Q3=6.00) and satisfaction with their control (treatment: M=5.00,Q1=2.75,Q3=6.00, baseline: M=5.00,Q1=3.00,Q3=5.25) under the treatment condition only if we look within-subjects (Figure 7a-b). Eleven participants had higher perceived control with Papers-to-Posts than the baseline tool (equal: 2, lower: 7), and 10 participants had higher satisfaction with Papers-to-Posts (equal: 4, lower: 6). On the other hand, participants did not demonstrate increased perceived ownership (treatment: M=6.00,Q1=5.75,Q3=6.00, baseline: M=6.00,Q1=5.00,Q3=6.00) or satisfaction with ownership (treatment: M=6,Q1=4.00,Q3=6.25, baseline: M=6.00,Q1=5.00,Q3=6.00) under the treatment condition (Figure 7c-d). Although they felt that they had more control over Papers-to-Posts, participants may not necessarily have felt that the LLM tool contributed more or less to the creation of the blog post.
5.5.6. Fulfillment of Design Goal 1
Looking at Figure 8a-b, we see that participants were more confident that all essential information was in the blog post (treatment: M=6.00,Q1=3.00,Q3=6.25, baseline: M=5.00,Q1=3.00,Q3=6.00) and found it easier to include essential information from the paper (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.75,Q3=6.00) using Papers-to-Posts. In addition, the interaction logs show that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post. All participants adjusted the LLM’s pre-selected bullet points to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs. We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. Participants also provided an average of 1.70 custom bullet points.
In their short answers regarding aspects of Papers-to-Posts that they found difficult and helpful, 10 participants noted the paper outline being useful in some manner. Four participants described the outline as helpful for generating text. For instance, P15 noted, ”I could rewrite entire sections or direct it with just a few bullets.” Three participants mentioned that the outline was useful for reviewing the information in or structure of the paper. P4 commented, ”The outline was really helpful in providing a quick overview of what content I had written.” For two participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P6 explained, ”The checked-marked bullet points helped me understand where the pre-filled content on the left was from,” and P12 shared, ”I liked the bullet point features to see what was being fed in vs not.” Five participants also specified a difficulty related to the paper outline. Two participants were overwhelmed with the number of bullet points. P9 described the outline as ”really bloated and hard to navigate.” Three more participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3 ”felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets,” and P20 did not realize that bullet points could come from different sections of the paper.
5.5.7. Fulfillment of Design Goal 2
Participants’ Likert responses regarding DG2 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) and iterate on the blog post text (treatment: M=6.00,Q1=5.00,Q3=6.00, baseline: M=5.00,Q1=4.00,Q3=6.00) in the treatment condition (Figure 8c-d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating and modifying text for sub-tasks (Appendix 27). For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and the context of prior blog post sections (6.00 times), though we note that these features were toggled on by default. In addition, they modified text 6.85 times on average, with the ”condense” (1.85 times), ”custom” (1.80 times), and ”simpler terms” (1.40 times) modifications being the most common.
In their short answers, 5 participants noted the usefulness of being able to provide custom instructions for either generating or modifying text using Papers-to-Posts. P9 described the custom instructions as providing a greater sense of control: ”…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” With respect to revising writing in particular, 12 participants noted in their short answers that Papers-to-Posts’s affordances for modifying text were useful. All of the preset modification buttons (”simpler terms”, ”condense”, ”expand”, ”more dramatic”, ”less dramatic”) were specifically mentioned as helpful affordances at least once. The most commonly noted one was the ”condense” button. For example, P12 commented, ”I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Two participants commented that they desired more preset modification buttons. P16 wanted a ”more technical” button, while P7 simply shared, ”I wish there are more default buttons for similar functions.” P7 then added, ”Or even let user to add their favorite rewrite button.” Interestingly, P2 found the preset modification buttons useful for understanding the LLM tool’s capabilities. They reflected, ”It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ’please summarize the following’.”
Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts. Eight participants commented that they appreciated the simplicity of the baseline tool or that the baseline tool gave them more freedom over the instructions sent to the tool. For instance, P9 noted, ”the ’generate writing’ panel felt overwhelming in terms of features,” and P19 explained, ”It’s very open end so I felt like I got more freedom what to ask it to do.” Meanwhile, in the treatment condition, 4 participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts. As an example, 2 participants wanted assistance with checking for redundancies in the blog post. Another 2 wanted feedback on the blog post writing. Regarding the affordances for modifying text in particular, 8 participants noted a relative difficulty. A common issue, raised by 4 participants, was confusion about how or when to use certain affordances. As an example, P3 ”didn’t really understand if the custom instructions only applied for the ’custom modification’ option,” when it in fact would be applied to any modification action. Meanwhile, P6 was unsure of ”when to use which one of the six [modification] options.” Another recurrent issue noted by 5 participants was a modification affordance not working as desired. For P20, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12, modifications were sometimes too exaggerated, while for P16, a custom modification to make writing more technical was too subdued. Despite these comments about feeling restricted or confused by Papers-to-Posts, participants still reported higher perceived control and satisfaction with control in the treatment condition, which we discuss in Section 5.5.5.
5.5.8. Extent of Change to Blog Post
Interestingly, we found that participants made larger changes to the initial LLM draft blog post when using Papers-to-Posts as compared to the baseline tool. In Figure 9a, for each condition, we show the Levenshtein distance, a common metric for measuring the amount of editing done to text, between participants’ initial LLM draft blog posts and their blog posts at the point at which the writer would feel comfortable publishing it publicly. For exploratory statistical testing, we employed a paired-samples t-test and observed that this Levenshtein distance is significantly higher with Papers-to-Posts (M=3977.05,SD=1146.46) compared to the baseline (M=2764.15,SD=1042.16) (Paired-Samples t-Test, t(19)=4.31, p¡.0005).
In Figure 9a, we show how the Levenshtein distance progressed over time, with the removal of one outlier. The difference in Levenshtein distance does not appear attributable to the time spent on the task or the change in length of the blog post. Participants did not spend substantially more time to reach the point at which they were comfortable publicly publishing the blog post using Papers-to-Posts (M=49.37, Q1=35.43, Q3=59.54) as compared to the baseline (M=46.81, Q1=33.92,Q3=56.60) (Appendix 25). Furthermore, we saw that participants changed the length of their blog posts, measured in characters, less using Papers-to-Posts (M=-162.00, Q1=-1126.25, Q3=297.00) than using the baseline (M=1971.00, Q1=965.25,Q3=2514.25) (Appendix 25(b)). Thus, we suspected that Papers-to-Posts may have required a decreased number of writing actions (inserting, deleting, replacing text in text fields) in order to achieve a given increase in Levenshtein distance. We investigated this potential reduction in effort in the follow-up deployment study.
6. Deployment Study
Through the controlled lab study, we observed that participants produced final blog posts that were significantly more distant from their initial LLM-generated drafts when using Papers-to-Posts than the baseline tool. We wanted to further understand how participants in a more realistic setting perceive Papers-to-Posts versus the baseline tool (RQ2), and whether Papers-to-Posts would lead participants to change their provided drafts more in a given time period (RQ1). For this, we conducted a deployment study where participants wrote blog posts on their own time with the explicit goal of posting the article publicly and sharing it with others.
6.1. Participants
@@
6.2. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors (Appendix 29). We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model with a larger context window that became available after the lab study. This enabled Papers-to-Poststo support longer inputs. In the deployment study, participants could freely close and reopen the tool in their browsers to work on their drafts when convenient. Participants could save their progress across sessions with a newly added save button. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long. Lastly, the treatment tool now automatically parsed the input paper.

### v2

@@ 4.1.1. Hypotheses
We hypothesized that, compared to a baseline consisting of an LLM draft and the ability to prompt the LLM, Papers-to-Posts would provide participants a better experience and outcome. Our individual hypotheses were as follows:
@@
H4: Papers-to-Posts leads to less time spent generating a blog post that the writer would be comfortable posting publicly, in comparison to the baseline.
4.1.2. Participants
@@ 4.1.3. Study Conditions
The treatment condition involved interacting with Papers-to-Posts which is described in Section 3. The tool in the baseline condition was designed to be a simplified version of Papers-to-Posts without affordances developed to address the design goals (Figure 4). The baseline provided input paper text, sourced from the PDF and divided into paragraphs and collapsible sections as in Papers-to-Posts. The baseline also provided the ability to prompt GPT-4 with any instructions and copy its output. A history of generations and their associated instructions was retained for the user’s perusal. Lastly, the baseline had an editor for writing the blog post, which was pre-populated with an LLM draft generated by GPT-4-32k using the entirety of the input paper.444An example of a baseline initial draft is in Appendix E. GPT-4-32k was used rather than GPT-4 because the context window for GPT-4 was insufficient for longer research papers. The prompt was designed to mirror the prompt for Papers-to-Posts’s initial draft sections (Appendix B.1), so the resulting draft also had introduction, methods, results, and conclusion sections. However, the prompt did not contain section-specific instructions.
@@
In each condition, the tool was set up for one of the participant’s research papers submitted in the recruitment survey. The paper was different for each condition. Each paper was required to have been published online between 2018 and the study period in 2023, with the participant as an author. Only one paper was published in 2018, and the rest were published during or after 2020. There was no significant difference in publication dates between conditions (Paired-Samples t-Test, t(19)=0.90, p=n.s.). The average paper length across conditions differed by less than 10% (baseline: 8452 words, treatment: 7742 words).
4.1.4. Study Procedure
@@
The study was within-subjects, so each participant completed two study sessions, one in the treatment condition and one in the baseline, in randomized order and following a counter-balanced design. The participants’ selected papers were also assigned to conditions in randomized order. Each study session lasted up to about 75 minutes. For sufficient flexibility, participants selected two study sessions within a week of each other, but the majority of sessions were a day apart (1 day: 13, 2 days: 3, 3 days: 3, 4 days: 1).555The condition order was counter-balanced within the group of participants with sessions one day apart and within the group of participants with sessions over one day apart.
@@
Participants completed a participation agreement before the study. We recorded and transcribed each session in Google Hangouts.666The lab study session script may be found in the supplementary materials. To start, the session coordinator introduced the task and presented a video tutorial on the assigned tool.777See the tutorial scripts in the supplementary materials. Participants then had two minutes to interact with and ask questions about the tool in the context of a sample paper that was not their own. Next, participants opened a tool link initialized with one of their research papers. The coordinator asked them to assume that their audience was people in their broad area (i.e., computer science), which was the most common audience for whom potential study recruits wanted to write blog posts (52/60). The coordinator instructed participants to work as if they were not there, but to let them know if any questions about the tool arose. They also instructed participants not to use any external tools or information. We made an exception twice when participants wanted to view their paper’s PDF due to issues with its parsing in the tool. Lastly, the coordinator asked participants to let them know if they reached a point at which they would feel comfortable publicly publishing the blog post in its current state.
@@
Each participant had up to one hour to work on the blog post. At the 20-minute and 40-minute marks, they were reminded to let the coordinator know if they reached the point of comfort with publication. If and when the participant reached that point, the coordinator asked them to save their draft and continue revising the blog post. After the first four study sessions, we noticed that participants were sometimes done early and did not desire to revise further. Thus, we added an additional rule that the participant would let the coordinator know if they reached a point at which they were proud of the blog post and saw no reason to revise further. If and when the participant reached that point, or if an hour had passed, they submitted their final blog post and moved on to a survey regarding their experience with the tool. The survey888The survey may be found in the supplementary materials. consisted of Likert-type questions regarding perceived efficiency, cognitive load, ease of using the tool in relation to the design goals, satisfaction with the tool, satisfaction with the final blog post, perceived control, and perceived ownership.999See Appendix A for a discussion of the exploratory results regarding participants’ perceived control and ownership. There were also short answer questions regarding the difficulties and benefits of using the tool in relation to the design goals.101010If it was the participant’s second session, the survey ended with a few questions directly comparing the two conditions. These were not for analysis but only to help us get a quick sense of the results. At the session’s end, the coordinator told the participant to close the tool and not to access it before the next session if they had one. Each time the participant prompted the LLM or submitted the blog post, a log of the action along with the state of the blog post was saved for analysis.
4.2. Deployment Study
We conducted a deployment study to understand if and how the results of the lab study translate to a less controlled and more realistic setting. To make the deployment study more realistic than the lab study, participants wrote blog posts on their own time with the explicit goal of sharing the blog post with others (e.g., on social media, in an email to colleagues).
4.2.1. Research Questions
Having observed that the lab study participants had significantly more editing power when using Papers-to-Posts rather than the baseline tool, we investigated if Papers-to-Posts leads participants to have greater editing power in this real-world context (RQ1). As in the lab study, we also evaluated satisfaction with the final blog post, satisfaction with the tool, and cognitive load (RQ2). We did not evaluate task completion time in the deployment study, as the lab study did not demonstrate any interesting difference between the conditions.
RQ1: How is potential users’ editing power impacted by Papers-to-Posts versus the baseline tool in a real-world setting?
RQ2: How do potential users perceive the outcome (blog post satisfaction (a)) and experience (tool satisfaction (b), cognitive load (c)) of interacting with Papers-to-Posts versus the baseline tool in a real-world setting?
For the deployment study, we posed research questions rather than evaluated hypotheses for statistical significance because power analyses indicated that we would need a larger sample size than we could obtain to see the same significant results from the lab study.111111The significant editing power result would have required 29 participants per condition (power=0.95, alpha=0.05, effect size=0.45), and the significant blog post satisfaction result would have required 83 participants per condition (power=0.95, alpha=0.05, effect size=0.98).
4.2.2. Participants
@@
4.2.3. Study Conditions
While the treatment and baseline conditions were largely the same as in the lab study, several minor usability issues identified in the lab study were addressed before proceeding with the deployment study. For example, there were better instructions and less intrusive highlight colors (Appendix 28). We also replaced GPT-4 and GPT-4-32k with the newer GPT-4-0125-preview model, which had a larger context window and became available after the lab study. This enabled Papers-to-Posts to support longer inputs. In the deployment study, participants could freely close and reopen their assigned tools in their browsers to work on their drafts when convenient. Participants could save their progress across sessions with a newly added save button. We note that, due to the manner in which text was logged in the deployment study, the tool in both conditions had a lag if participants pasted a chunk of text more then a couple paragraphs long. Lastly, the treatment tool now automatically parsed the input paper.